OpenStudy (calculusxy):

Write the interval notation of the compound inequality 6b - 1 \(\le\) 41 or 2b + 1 \(\ge\) 11

OpenStudy (calculusxy):

@SolomonZelman

OpenStudy (calculusxy):

@jhonyy9 @KamiBug @zepdrix

jhonyy9 (jhonyy9):

6b-1<= 41 what will get than add to both sides 1 ?

OpenStudy (calculusxy):

b \(\le\) 7

OpenStudy (calculusxy):

and for the other one it's b \(\ge\) 5

jhonyy9 (jhonyy9):

yes this is after divide both sides by 6

OpenStudy (calculusxy):

When I put a number line, I saw that there was one section that shared between both the inequalities. It was between 5 and 7.

OpenStudy (calculusxy):